What would be the pH of a solution in which the molar ratio of propanoic acid: sodium propanoate is 1:10 ka for propanoic acid is 1.34×10^-5 mol/dm^3

1
Expert's answer
2021-11-09T10:28:34-0500

"[H^+_{(aq)}]=\frac{Ka[Acid_{(aq)}]}{salt_{(aq)}}"

"=1.34×10^{-5}×\frac{1}{10}"

"=1.33\u00d710^{-6}" mol dm-3


pH= "-log[H^+_{(aq)}]=- log (1.34\u00d710^{-6})"

"=5.87"